OpenStudy (amy2012):

A total of 27000 is invested, part of it at 12% the rest at 13%. the total interest after one year is 3385 how much is invested at each rate

OpenStudy (phi):

part of it at 12% the rest at 13%. how many parts is that ? I think 2. give the two parts "names" like x and y any idea what x and y add up to ?

OpenStudy (amruta1202):

Chk this out

OpenStudy (amruta1202):

Where I1 and I2 are interests at 12% and 13% respectively where I= p*r*n/100

OpenStudy (dumbcow):

set up 2 equations Money invested: x + y = 27000 Interest: .12x + .13y = 3385 solve for x,y using substitution ---> x = 27000 - y sub into 2nd equation